what is the probability of rolling a number greater than 1 on a regular six-sided die?
Answer #1

Six sided die is S={ 1, 2, 3, 4, 5, 6 }

Total event = 6

Roll die probability
1 1/6 0.1667
2 1/6 0.1667
3 1/6 0.1667
4 1/6 0.1667
5 1/6 0.1667
6 1/6 0.1667

Now, we have to find P(X > 1)

Therefore,

P(X>1) = P(X=2) + P(X=3) + P(X=4) + P(X=5) + P(X=6)

P(X>1) = 1/6 + 1/6 + 1/6 + 1/6 + 1/6

P(X>1) = 5/6

P(X>1) = 0.8333